40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Oct 8 18:35:09 2019, gautam, Update, LSC, Locking prep 
    Reply  Wed Oct 9 18:15:26 2019, gautam, Update, LSC, PRMI 3f locking PRMI_1f.png
       Reply  Wed Oct 9 22:02:58 2019, gautam, Update, LSC, REFL55 whitening issue 
          Reply  Thu Oct 10 01:12:56 2019, gautam, Update, LSC, Locking studies PRFPMIstudies.png
             Reply  Wed Oct 16 11:42:17 2019, gautam, Update, LSC, Poor separation of PRCL/MICH in 3f signals sensMat.pdf
                Reply  Thu Oct 17 11:19:28 2019, gautam, Update, LSC, Locking activity last night MICH2PRCLnulling.pdfconsolidatedSensingMatrices.pdf.zip
                   Reply  Tue Oct 22 00:52:27 2019, gautam, Update, LSC, Locking updates PRMI3f_ALS_Oct21sensMat.pdf
                      Reply  Tue Oct 29 01:24:45 2019, gautam, Update, LSC, More locking updates POP22_feature.pngPRMI3f_ALS_Oct21sensMat.pdf
                         Reply  Tue Oct 29 15:13:19 2019, gautam, Update, LSC, More locking updates PRMI_lockloss.png
                            Reply  Wed Oct 30 01:27:00 2019, gautam, Update, LSC, More locking updates POP_ASS.png
                               Reply  Wed Oct 30 11:53:41 2019, gautam, Update, LSC, MICH loop shape tuning MICH_ELP.pdfMICH_ELP_TFs.pdf
                               Reply  Wed Oct 30 17:08:40 2019, gautam, Update, LSC, POP22 investigation POP22anomaly.pdf
                                  Reply  Mon Nov 4 15:29:47 2019, gautam, Update, LSC, POP signal path POPchain.pdf
                                     Reply  Mon Nov 4 16:06:58 2019, gautam, Update, LSC, POP optical path IMG_8027.JPGlensRealignment.jpgPOPrealigned.pngPOPdither.pngPRMfixed.png
                                        Reply  Tue Nov 5 11:52:27 2019, gautam, Update, LSC, Locking notes PRMI3f_ALS_Nov4sensMat.pdf
                                           Reply  Wed Nov 6 02:08:48 2019, gautam, Update, LSC, Locking updates 
                                              Reply  Wed Nov 6 17:05:45 2019, gautam, Update, LSC, CARM calibration CARMscan.pdfCARMcalib.pdfscan.hdf5.zip
                                              Reply  Wed Nov 6 17:45:34 2019, gautam, Update, LSC, ~ 
                                                 Reply  Mon Nov 18 21:04:38 2019, gautam, Update, LSC, Locking - some ideas 
                                                    Reply  Wed Nov 20 01:07:18 2019, gautam, Update, LSC, Locking - progress PRFPMI.png
                                                       Reply  Wed Nov 20 12:14:17 2019, gautam, Update, LSC, Locking - progress TRX_TRY_comparison.pdf
                                                          Reply  Wed Nov 20 17:20:24 2019, Yehonathan, Update, LSC, QPD Investigation 
                                                             Reply  Wed Nov 20 17:52:00 2019, gautam, Update, LSC, QPD MEDM screen update newLookQPD.pngTRX_TRY_comparison.pdfIMG_8186.JPG
                                                                Reply  Wed Nov 20 21:29:28 2019, gautam, Update, LSC, PRG ~13 PRG13.pdf
                                                                   Reply  Fri Nov 22 00:54:14 2019, gautam, Update, LSC, locking notes ALSnoiseIncrease.pdf
                                                                      Reply  Mon Nov 25 19:11:22 2019, gautam, Update, LSC, ALS noise re-look ALSnoise.pdf
                                                                         Reply  Wed Nov 27 23:24:01 2019, gautam, Update, LSC, No shaking but no inspiration either 
                                                                            Reply  Sun Dec 1 13:38:48 2019, rana, Update, LSC, No shaking but no inspiration either 
                                                                               Reply  Tue Dec 3 00:03:57 2019, gautam, Update, LSC, Green ALS also shows elevated noise with high arm buildup ALSnoiseIncrease_greenBeat.pdf
                                                                                  Reply  Wed Dec 4 19:54:27 2019, gautam, Update, LSC, A look to the past DRFPMI_2016March.pdf
                                                                                     Reply  Thu Dec 5 01:54:39 2019, gautam, Update, LSC, Partial CM board path engaged AOengaged.pdf
                                                                                        Reply  Fri Dec 6 00:02:48 2019, gautam, Update, LSC, What is the correct way to set the 3f offsets? 
                                     Reply  Fri Nov 15 12:08:04 2019, gautam, Update, LSC, POPDC whitening board 
                         Reply  Tue Oct 29 17:40:48 2019, gautam, Update, LSC, More locking updates broadSweep.pdfzoomSweep.pdf
Message ID: 14998     Entry time: Tue Oct 29 17:40:48 2019     In reply to: 14996
Author: gautam 
Type: Update 
Category: LSC 
Subject: More locking updates 

I set up a photodiode (PDA10CF) in the IFO REFL beampath and the Agilent NA is sitting on the east side of the PSL enclosure. This was meant to be just a first look, maybe the PDA10CF isn't suitable for this measurement. The measurement condition was - PRM aligned so we have a REFL beam (DC level = 8.4V measured with High-Z). Both ITMs and ETMs were macroscopically misaligned so that there isn't any cavity effects to consider. I collected noise around 11 and 55 MHz, and also a dark measurement, plots to follow. The optics were re-aligned to the nominal config but I left the NA on the east side of the PSL enclosure for now, in anticipation of us maybe wanting to tune something while minimizing a peak.

Attachment #1: Results of a coarse sweep from 5 MHz to 100 MHz. The broadband RIN level is not resolvable above the dark noise of the photodiode, but the peaks at the modulation frequencies (11 MHz, 55 MHz and 29.5 MHz) are clearly visible. Not sure what is the peak at ~44 MHz or 66 MHz. Come to think of it, why is the 29.5 MHz peak so prominent? The IMC cavity pole is ~4kHz so shouldn't the 29.5 MHz be attenuated by 80dB in transmission through the cavity?

Attachment #2: Zoomed in spectra with finer IF bandwidth around the RF modualtion frequencies. From this first measurement, it seems like the RIN/rad level is ~10^5, which I vaguely remember from discussions being the level which is best achieved in practise in the 40m in the past.

Quote:
 

Check the RAM due to the EOM? Perhaps the pointing / polarization control into the EOM got degraded.

Attachment 1: broadSweep.pdf  184 kB  Uploaded Wed Oct 30 01:22:54 2019  | Show | Hide all | Show all
Attachment 2: zoomSweep.pdf  191 kB  Uploaded Wed Oct 30 01:23:00 2019  | Hide | Hide all | Show all
zoomSweep.pdf
ELOG V3.1.3-